OpenStudy (coplops):

A sequence is defined by the recursive function f(n + 1) = f(n). If f(3) = 9 , what is f(1) ? 1 3 27 81

OpenStudy (holsteremission):

If \(f(n)=f(n+1)\), then every term in the sequence is the same.

OpenStudy (coplops):

So how would I get the answer?

OpenStudy (queelius):

As Holster indicated, if f(n+1)=f(n), then f is a constant function. We are given the value of this function, i.e., f(3) = 9. Therefore, f(n) = 9. So, f(1) = 9 also. Unfortunately, this is not an option among the multiple choices. This leads me to suspect you didn't present the problem correctly.